OpenStudy (anonymous):

Solve for the variable: g/24=-2

jimthompson5910 (jim_thompson5910):

You're dividing by 24, so you need to undo this and do what?

Parth (parthkohli):

Multiply both sides by 24. What do you get?

Parth (parthkohli):

@jim_thompson5910 :) Isolation of the variable, exactly.

OpenStudy (anonymous):

honestly dont know

jimthompson5910 (jim_thompson5910):

Start with g/24 = -2 and multiply both sides by 24

jimthompson5910 (jim_thompson5910):

so 24*(g/24) is 24g/24 which becomes g

jimthompson5910 (jim_thompson5910):

and -2*24 = -48 So g/24 =-2 turns into g = -48

OpenStudy (anonymous):

ok i get it now can you help me on a new question ? based on the same thing .. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure whats on your mind?

OpenStudy (anonymous):

let me close this question and retype it

jimthompson5910 (jim_thompson5910):

alright